OpenStudy (anonymous):

d/dx ln[x(x^2+3)^3] = ? How to solve this?

OpenStudy (anonymous):

\[\frac{ d }{ dx } \ln \left[x \left( x^2+3 \right)^3 \right]\] is this the function?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so\[\frac{ d }{ dx}\ln u = \frac{u'}{ u }\]

OpenStudy (anonymous):

you can multiply or use product rule for x(x^2+3)^3

OpenStudy (anonymous):

I already tried but couldn't get the right result

OpenStudy (anonymous):

\[\frac{(x ^{2}+3)^{3} +3(x^{2}+3)^{2}2x}{x(x^{2}+3)^{3}}\]

OpenStudy (anonymous):

is this right?

OpenStudy (anonymous):

no. \[u = x(x^2+3)^3 \Rightarrow u' = (x^2+3)^3+x(2)(x^2+3)^2(2x)=(x^2+3)^2(x^2+3+4x^2)\] \[=(5x^2+3)(x^2+3)^2\]

OpenStudy (psymon):

Hmm.... \[(x^{2}+3)^{3} + x(3)(x^{2}+3)^{2}(2x) = (x^{2}+3)^{2}[(x^{2}+3)^{1}+ 6x^{2}] = (x^{2}+3)^{2}(7x^{2}+3)\] \[\frac{ (x^{2}+3)^{2}(7x^{2}+3) }{ x(x^{2}+3)^{3} }\implies \frac{ (7x^{2}+3) }{ x(x^{2}+3) }\]

OpenStudy (psymon):

Aww, went off the edge

OpenStudy (anonymous):

oops, should be 3x^2 giving just what @psymon wrote.

OpenStudy (anonymous):

ah thank you, I think the answer of this question is wrong, it is (1-x^2)/x(x^2+1), which is not related at all =.=

OpenStudy (psymon):

lol.

OpenStudy (anonymous):

where did that come from?

OpenStudy (anonymous):

my homework from my professor

OpenStudy (anonymous):

the solution? perhaps your prof was looking at a different prob...

OpenStudy (anonymous):

yep, i think so, he sometimes makes mistakes at class xD

OpenStudy (anonymous):

it happens... it's so difficult to be perfect! in fact, i'd say impossible.